OpenStudy (anonymous):

3x+(5/2)y=21 x+2y=14 can someone help me solve this plzzz.

OpenStudy (kropot72):

You can find x in terms of y by rearranging the second equation as follows: x + 2y = 14 Therefore: x = 14 - 2y Do you understand this first step?

OpenStudy (anonymous):

your subtracting 2y from both sides?

OpenStudy (kropot72):

Yes, you are correct. Now we can plug this expression for x into the first equation 3x + (5/2)y = 21 giving 3(14 - 2y) + (5/2)y = 21 which expands to 42 - 6y + (5/2)y = 21 Can you now solve this last equation to find the value of y?

OpenStudy (anonymous):

first I think is 21 - 42 which is -2. 6y + (5/2)y= -21 after that i'm kinda confused

OpenStudy (usukidoll):

\[3x+ \frac{5}{2}y=21\] \[x+2y=14\] I think dealing with the fraction is a chore... maybe we can multiply 2 throughout the first equation so we can have whole numbers to work with

OpenStudy (anonymous):

yea im not good with fractions

OpenStudy (usukidoll):

\[3(2)x+ \frac{5(2)}{2}y=21(2)\] \[6x+ 5y=42\] \[x+2y=14\]

OpenStudy (usukidoll):

we could use subsitution by using the second equation x=14-2y and then plug it into the first equation

OpenStudy (usukidoll):

\[6(14-2y)+ 5y=42\]

OpenStudy (kropot72):

The equation simplifies to - (7/2)y = -21 Now you can multiply both sides by -2 to give 7y = 42

OpenStudy (anonymous):

6

OpenStudy (kropot72):

yes y = 6 is correct.

OpenStudy (usukidoll):

\[84-12y+5y=42 \rightarrow 84-7y=42\] \[-7y=42-84 \rightarrow -7y=-42 \rightarrow y=6\]

OpenStudy (usukidoll):

yup y = 6

OpenStudy (anonymous):

yay :D

OpenStudy (usukidoll):

so now we can plug y = 6 back into one of the equations.. let's try the second equation x+2y = 14

OpenStudy (usukidoll):

x+2(6)=14 x+12=14

OpenStudy (kropot72):

Which shows there is more than one way to skin a cat, if you forgive the analogy :)

OpenStudy (usukidoll):

ewwww I own cats in real life

OpenStudy (anonymous):

x=2 n I had a cat once

OpenStudy (kropot72):

My bad form :(

OpenStudy (usukidoll):

yes so x = 2 and y = 6 we can plug it back into the original equations to see if we have the right answer.. I know we do because 2+2(6)=14 2+12=14 14=14

OpenStudy (anonymous):

thank u

OpenStudy (usukidoll):

3(2) +5/2(6/1) = 21 6+(30/2)=21 6+15=21 21=21

OpenStudy (anonymous):